# Knowledge-Centric Templatic Views of Documents

### v1

@@ 5.1 Quality Measure
For the TAE precision score, we calculate the average similarity between the generated panels and their most similar reference panel. More specifically, we calculate $Q_{P}$ as follows:
@@
in which $S$ is the set of reference panels and $\tilde{S}$ is the set of generated panels. For the similarity metric, the user can choose a similarity metric that best matches their use case, such as ROUGE, BERT-Score, or a custom trained regressor Lin (2004); Zhang* et al. (2020). For example, a user might choose ROUGE if they want a similarity metric that focuses on exact word overlap, or they may choose BERTScore to measure broader semantic similarity.
Similar to precision, for the TAE recall score, we calculate the average similarity between the reference panels and their most similar generated panel:
@@
In other words, we measure how similar each reference panel is to the generated panels.
@@ 5.2 Order Penalty
Formally, assume $S$ and $\tilde{S}$ are sequences of reference and generated panels respectively. We use the maximum similarity scores calculated in §5.1 to align the panels across sets. For the precision ordering penalty, we define the following operation $\lambda_{P}(s)=\sum_{\tilde{s}}\delta_{P}(s,\tilde{s})$, where
@@
Intuitively, this captures the cardinality of the alignment of a given panel in $S$ with panels in $\tilde{S}$. Then, using this operation we can replace every panel $s\in S$ with $\lambda(s)$ copies, leading to an identical cardinality for both $S$ and $\tilde{S}$, and subsequent one-to-one mapping between their corresponding panels.
With this one-to-one mapping in place we can then measure the misalignments between the two sets of ordered panels to assign a penalty scores. To operationalize this, we associate panels with ranks in both sets and use a rank correlation metric to compute the degree of agreement. Specifically, rank assignment is done as follows: panels in $\tilde{S}$ are simply assigned ranks in order of appearance 1 through N – we call this $\tilde{S}_{ranking}^{P}$; meanwhile panels in $S$ are assigned the identical rank to their one-to-one aligned panel in $\tilde{S}$ and $\tilde{S}_{ranking}^{P}$ – we refer to these rankings as $S_{ranking}^{P}$. The final ordering penalty is computed using Spearman’s rank correlation Szmidt and Kacprzyk (2010):
@@
where we perform a linear transformation to map the original range of the correlation coefficient [-1, 1] to the desired range [0, 1] for precision.
Similarly, for the recall ordering penalty, we perform the same operation but in the other direction, mapping the reference panels to the generated panels. This means that panels st $\delta_{P}(s,\tilde{s})=0$ will be accounted for in the recall ordering penalty. Formally, $\lambda_{R}(s)=\sum_{s}\delta_{R}(\tilde{s},s)$, where
$\delta_{R}(\tilde{s},s)=\begin{cases}1,&\text{iff }\tilde{s}\rightarrow s\\ 0,&\text{otherwise}\end{cases}$
We calculate the rankings for recall similarly, assigning the panels in $S$ a rank of $S_{ranking}^{R}=[1,...,N]$ in order of appearance, while panels in $\tilde{S}$ are assigned the identical rank to their one-to-one aligned panel in $S$ and $S_{ranking}^{R}$. We then take the Spearman’s rank correlation as follows:
$O_{R}=\frac{\mathrm{Spearman}(S_{ranking}^{R},\tilde{S}_{ranking}^{R})+1}{2}$ (5)
An example of this process can be found in Figure 3.
Figure 3: Example of the process of obtaining the rankings for the precision ordering penalty. This process is reflexive, and panels not accounted for in the precision ordering penalty are accounted for in the recall ordering penalty.
Simularity Measure
Template Model Template R-1 R-L BERTScore
slides GPT4 SURe 8.95 8.37 38.08
Base 8.89 8.29 36.24
[…]
Base 5.62 5.09 31.64
[…]
Base 27.73 25.68 82.76
[…]
Base 27.02 25.10 82.85
[…]
Base 17.61 16.35 59.82
[…]
Base 8.38 7.55 35.45
Table 2: Evaluation results using GPT3.5 (gpt35-16k) and GPT4 (gpt4-32k). For each model, we compare the performance with and without the SURe. We report the TAE F1 scores as calculated in §5, using Rouge-1 (R-1), Rouge-L (R-L), and Bert-Score as the similarity metrics.
@@ 5.3 Length Penalty
If $L$ were reflexive, the denominator in the reverse case would be $|\tilde{S}|$, meaning that as $|\tilde{S}|\rightarrow\infty$, $L\rightarrow 1$. In other words, the length penalty could be cheated by over-generating. Therefore, unlike the quality measure and ordering penalty, we chose not to make the length penalty reflexive.
@@ 6 Results
As a baseline, we skip the SURe generation step, passing the full text of the paper as input to the generate a LaTeX document prompt, rather than the SURe. We maintain the same style parameters for both settings. Past work on the Paper-Poster dataset and the DOC2PPT dataset did not release their code or results, so we cannot directly compare to their work. LongSumm’s test set is a blind set that does not allow us to compare using our custom metric.
We use gpt4-32k and gpt35-16k for our experiments. We truncated inputs that are too long for the input window and used a temperature of 0.0, based on hyper-parameter experiments described in §B.
@@ 6.1 Results of automatic evaluation
We report the results of our automatic evaluation in Table 2, computed the TAE scores as described in §5. As we can see from the results, first generating the SURe results in performance gains for most settings. We see that largest performance gains in the gpt35-16k experiments, indicating that our method has the largest benefits for simpler models. This is likely because more simple models have more difficulty parsing long unstructured documents, and the SURe provides a structured, filtered representation of the input document. The only setting we don’t see performance gains or similar performance is in the gpt4-32k setting on the poster template. Upon closer inspection, the references in this dataset are very verbose, averaging 391 tokens total. Our method produces generally less verbose posters, averaging 265 total tokens compared to 345 tokens produced by the baseline. We hypothesize that by editing the style parameters to include information about verbosity and length, we can improve performance on posters and leave exploration of this to future work.
6.1.1 Importance of Structure
Our method provides a template JSON to the LLM to generate the SURe, before passing the resulting JSON back to the LLM to generate the final document. A natural question is, how important is the JSON format? In other words, do we see performance gains as a result of the filtering of information, or is the structure provided by the JSON important as well? In order to test this question, we conduct an additional experiment in which we take the text within the SURe, discard the JSON structure, and pass the resulting body of text to the LLM to generate the final document. We randomly sample 100 documents each from LongSumm and Doc2PPT for the blog and poster generation tasks, respectively. We use the entirety of the Paper-Poster dataset, since it contains less than 100 examples. We use gpt35-16k for this experiment. The results of this experiment can be found in Table 3. As we can see from the results, passing the text of the SURe without the JSON structure decreases performance in most cases, with posters having the largest gap in performance, indicating that both the structure and content of the SURe are important for performance.
Template
SURe
Format
R-1 R-L
Bert
Score
slides JSON 7.7 7.3 35.1
Text 7.9 7.1 34.6
@@ 6.1.1 Importance of Structure
Text 11.5 10.5 46.3
Table 3: TAE scores comparing the use of the JSON representation of the SURe or passing the text generated in the SURe, using gpt35-16k. We see that in most cases, the JSON representation performs better, indicating the importantance of structure.
Figure 4: Reasons annotators preferred each document. While annotators largely preferred documents generated with a SURe, the most common reasons for preference are better formatting and information content. We exclude the “Other” count as it was only selected once.
@@ 6.2 Human evaluation
We additionally conduct a human evaluation to test both our method and how well our evaluation framework aligns with human preference. We sample 100 documents each from Doc2PPT and LongSumm, and use the entirety of the Paper-Poster dataset. We present the annotators with 2 versions of each document, one generated with the SURe and one without. Both versions use gpt4-32k. The annotators are provided with the original paper and the intended document type (blog, slide deck, or poster), and are asked the following questions:
  1. 1.
    Which document do you prefer?
  2. 2.
    On a scale of 1-3, to what degree do you prefer your selection?
  3. 3.
    Why do you prefer your selection?
For Question 3, the annotators are given a multi-select checklist of reasons why they might prefer one document over the other. The checklist included (1) the quality of the content, (2) the formatting, (3) the document style matching the intended document types, (4) the information represented in the document, and (5) other. The “other” option included a free text box. The full annotation instructions, including a description of the reasons provided and examples, can be found in §A.
Each document was annotated by 3 different annotators. We employed 4 annotators from India, sourced via a third-party agency, to carry out the human evaluation of our task. A guideline document was provided to these evaluators, containing task-specific instructions, guidance, and several annotated examples. They were compensated at a rate of \$11.98 USD per hour for the total time that they spent working on the task, including a training round of annotation.
@@
The documents generated with a SURe were preferred by 82%, based on majority vote. 71% of the documents generated with a SURe were unanimously preferred by the annotators. We calculate a annotator agreement score of 0.51 with Krippendorff’s alpha, indicating that while this is a subjective and specialized task, even non-expert annotators agree to a moderate degree. A visualization of the reasons the annotators preferred their selection can be found in Figure 4. We can see from the results that while annotators largely preferred the documents generated with a SURe, the most common reasons for preference are better formatting and better information content. This indicates that the structure provided by the SURe makes it easier for the model to format the final document well. Additionally, the SURe only includes the most salient information from the original text, resulting in higher quality of information content. Finally, we see a fairly even distribution across the different templatic views for the reasons of preference, indicating humans prefer the documents generated with the SURe across different document types.
@@
BERTScore 10.6*
Table 4: Correlation of evaluation metrics with human judgement. We compare each metric computed using the TAE framework versus the standard computation. *Indicates the correlation is not statistically significant ($p>0.01$)
Which metric correlates better with human judgement?
We use the human preference judgements to test if our metric, as described in §5, correlates more with human preference compared to prior evaluation metrics in the literature. For each annotation, given the degree of the preference $d$ we convert the model preference to a score $P$ as follows:
$P(d)=\begin{cases}3,&d=\textrm{strong}\\ 2,&d=\textrm{moderate}\\ 1,&d=\textrm{slight}\end{cases}$
If the annotator prefers the document generated without a SURe, we multiply $P(d)$ by $-1$. This allows us to measure if the metric captures both directionality of preference and degree.
Finally, we compute the automatic score for each metric, then calculate $S=m(\textrm{with SURe})-m(\textrm{skip SURe})$, where $m$ is the metric we are evaluating. If the annotator prefers the document generated without the SURe, we’d expect a good metric to assign a higher score to that document as well, resulting in both $S$ and $P$ being negative. Conversely, we’d expect instances where the opposite were true to result in positive scores for both $S$ and $P$. Using these intuitions we assign an affinity score of a metric with respect to human evaluation as the Pearson correlation Freedman et al. (2007) of $S$ and $P$.
Since prior metrics are not designed to account for the structure of the document, we compute them by extracting only the text of both the generated document and reference documents. The correlations with human judgement for each metric, computed standard compared to TAE variants can be found in Table 4. As we can see from the results, computing each evaluation metric using our template agnostic evaluation method correlates more highly with human judgement, except in the case of BERTScore. However, BERTScore is the only metric in which the correlation is not statistically significant, as it is likely simply a poor metric for this task. We hypothesize that BERTScore does not perform well on scientific data, since it is trained primarily on out of domain text.
@@ 7 Conclusion
In many domains, people choose to disseminate information across different modalities and formats, allowing for better communication of information to broader audiences. Past work has focused on developing methods to generate different templatic views of documents, tailoring the method to each template. In this paper, we introduced an LLM-powered method to unify the generation of templatic views of documents by first generating a Structured Unified Representation (SURe) of the input document, representing the underlying knowledge of the templatic views. We showed that our method results in improvements in performance, especially for a simpler model. We additionally introduced Template Agnostic Evaluation (TAE), a flexible precision-recall framework for automatic evaluation that is template agnostic, allowing for easy integration of existing evaluation metrics into a unified system. The framework is capable of jointly accounting for and measuring multiple facets of the output document including the quality of its content, its underlying structure and ordering, as well as its length. Finally, we conducted a human evaluation and showed that annotators prefer the documents generated using the SURe 82% of the time. We additionally showed that our framework of evaluation correlates more highly with human preference than standard evaluation metrics.
The use of templatic views of documents exists in many domains (e.g. creating a cover letter and personal website based on a resume), and in future work we will consider other domains of documents, as well as more complex scenarios, such as many-to-many document generation. Our work is a step towards a larger goal of generalized document generation.
Limitations
Although our methods are not domain specific, we only evaluated our methods in the scientific domain, due to the availability of public data. Further research is required to evaluate the performance of our methods in other domains. Additionally, our method is limited to textual content. Future work should explore incorporating multi-modal models and content generation. Finally, we evaluated our methods on closed source models, which may affect the reproducibility of our results.
Appendix A Annotation Instructions
A.1 Questions

### v2

@@ 5.1 Quality Measure
For the TAE precision score, we calculate the average similarity between the generated panels and their most similar reference panel as follows:
@@
in which $S$ is the set of reference panels and $\tilde{S}$ is the set of generated panels. For the similarity metric, the user can choose a metric that best matches their use case, such as ROUGE, BERT-Score, or a custom trained regressor Lin (2004); Zhang* et al. (2020). For example, a user might choose ROUGE if they want a similarity metric that focuses on exact word overlap, or BERTScore to measure broader semantic similarity.
Similar to precision, the TAE recall score is calculated as the average similarity between the reference panels and their most similar generated panel:
@@ 5.2 Order Penalty
Formally, assume $S$ and $\tilde{S}$ are sequences of reference and generated panels respectively. We use the maximum similarity scores calculated in §5.1 to align the panels across sets.
For the precision ordering penalty, we define the following operation $\lambda_{P}(s)=\sum_{\tilde{s}}\delta_{P}(s,\tilde{s})$, where
@@
Intuitively, this captures the cardinality of the alignment of a panel in $S$ with panels in $\tilde{S}$. Then, using this operation we can replace every $s\in S$ with $\lambda_{P}(s)$ copies, leading to an identical cardinality for both $S$ and $\tilde{S}$, and subsequent one-to-one mapping between their corresponding panels.
Figure 2: Example of the process of obtaining the rankings for the precision ordering penalty. We first use the similarity measure to map each generated panel to its most similar reference document. This mapping is used to calculate the precision quality score $Q_{P}$. We then use the mappings to create a one-to-one alignment from the generated to the reference panels, which we use to calculate the precision ordering penalty ($O_{P}$). By creating a one-to-one alignment, we are able to represent inversions in the ordering. This process is reflexive, and panels not accounted for in the precision ordering penalty are accounted for in the recall ordering penalty.
Then, to operationalize a penalty score for the two sets of ordered panels we associate them with ranks in both sets and use a rank correlation metric to compute the degree of agreement. Specifically, rank assignment is done as follows: panels in $\tilde{S}$ are simply assigned ranks in order of appearance 1 through N – we call this $\tilde{S}_{ranking}^{P}$; meanwhile panels in $S$ are assigned the identical rank to their one-to-one aligned panel in $\tilde{S}$ and $\tilde{S}_{ranking}^{P}$ – we refer to these rankings as $S_{ranking}^{P}$. An example of this process can be found in Figure 2. The final ordering penalty is computed using Spearman’s rank correlation Szmidt and Kacprzyk (2010):
@@
where we perform a linear transformation to map the original range of the correlation coefficient [-1, 1] to the desired range [0, 1].
Similarly, for the recall ordering penalty, we map the reference panels to the generated panels, calculated as $\lambda_{R}(s)=\sum_{s}\delta_{R}(\tilde{s},s)$. $O_{R}$ is calculated similar to $O_{P}$, using the recall rankings.
@@ 5.3 Length Penalty
We chose to keep $L$ non-reflexive, because in the reverse case – as $|\tilde{S}|\rightarrow\infty$, $L\rightarrow 1$ – the metric could be cheated by over-generating.
@@ 6 Results
As mentioned in §3, past work on Doc2PPT and Paper-Posters do not release code, making it difficult to do a direct comparison. They also do not report any baselines to compare against. Meanwhile, Longsumm’s blind test does not allow us to compute our custom metric, although we do report the leaderboard results in Appendix D. Notably, with almost no prompt engineering our LLM-based system places second on this leaderboard. We argue that for the investigation in this paper, direct comparison to prior non-LLM baselines is not only unfair to those approaches, but not particularly insightful. Therefore, similar to Wei et al. (2022b), we focus on variants of our LLM-based method and treat them as baselines. Example outputs of each template type can be found in Appendix E.
We conduct experiments with the following settings: (1) No Representation – this is the default setting of going directly from the source document to the target document. We skip the intermediate generation step, passing the full paper as input. We experiment both with and without the style parameters. (2) Own Representation – we do not pass a JSON structure to the intermediate generation step, and allow the model to choose its own structure. (3) Text Representation – we extract the text from the intermediate representation, discarding the JSON structure. (4) JSON Representation – this is the full JSON structure for the intermediate generation step. We experiment both with and without the style parameters.
We use gpt35-16k in our main set of experiments. We truncate text that is too long for the input window and use a temperature of 0.0 as standard. [^5] A detailed evaluation of the temperature hyper-parameter is included in Appendix §C
Similarity Measure
Rep. Style R-L M B BERTS
Slides None $\times$ 5.0 6.4 0.3 31.6
None $\checkmark$ 5.1 6.0 0.4 31.7
Own $\checkmark$ 6.5 7.1 1.2 36.1
Text $\checkmark$ 7.3 8.0 1.4 36.4
JSON $\times$ 4.2 6.0 0.3 31.4
JSON $\checkmark$ 7.4 8.4 1.5 36.9
[…]
None $\checkmark$ 25.1 17.7 2.3 82.8
Own $\checkmark$ 23.9 19.2 2.3 82.2
Text $\checkmark$ 25.4 19.3 2.5 82.5
JSON $\times$ 28.3 25.3 5.0 82.3
JSON $\checkmark$ 25.4 19.6 2.8 82.4
[…]
None $\checkmark$ 10.1 11.6 1.9 39.5
Own $\checkmark$ 12.8 12.6 2.9 52.8
Text $\checkmark$ 11.3 11.7 2.1 45.9
JSON $\times$ 14.2 16.8 4.0 52.8
JSON $\checkmark$ 15.5 14.5 15.3 53.3
Table 1: Evaluation results using GPT3.5 (gpt35-16k). For each template, we experiment with different representations (Rep) and whether or not we include the style parameters (Style). We report the TAE F1 scores as calculated in §5, using ROUGE-L (R-L), METEOR (M), BLEU (B), and BERTScore (BERTS) as the similarity metrics.
@@ 6.1 Results of automatic evaluation
In Table 1, we report the TAE F1 scores as described in §5, using ROUGE-L Lin (2004), METEOR Banerjee and Lavie (2005), BLEU Papineni et al. (2002), and BERTScore Zhang* et al. (2020) for the similarity measure. As seen in the results, by most measures, generating a JSON intermediate representation yields the best performance.
We see that using the text representation generally degrades the performance over providing the structured JSON representation, indicating that structure is important for downstream performance in addition to abstractive filtering of information. Additionally, the text representation performs better than skipping the intermediate step altogether for both the poster and slide generation task, but not the blog generation task. This is likely because posters and slides have more inherent structure than blog posts, which can be relatively free-form.
Finally, we see that allowing the model to choose its own representation format degrades performance over providing our pre-defined JSON structure. However, we see that in most cases, providing a representation generated without a JSON structure still performs better than skipping the intermediate generation step altogether (while maintaining the same style parameter setting). This indicates that even without a pre-defined structure, the intermediate step is still valuable for performance.
Similarity Measure
Model Rep. R-L M B BERTS
Slides MS $\times$ 0.6 0.4 0.0 28.6
$\checkmark$ 4.4 4.6 0.4 30.5
@@
$\checkmark$ 8.4 9.1 2.0 38.1
[…]
$\checkmark$ 21.7 16.2 1.7 81.3
[…]
$\checkmark$ 25.6 20.5 3.2 82.5
[…]
$\checkmark$ 25.8 20.2 3.1 82.7
[…]
$\checkmark$ 6.0 6.5 1.2 38.1
[…]
$\checkmark$ 10.4 11.0 1.5 50.7
[…]
$\checkmark$ 14.6 15.3 3.7 57.2
Table 2: TAE F1 scores using Mistral-7b (MS), Mixtral (MX), and GPT4. We use ROUGE-L (R-L), METEOR (M), BLEU (B) and BERTScore (BERTS) as our similarity measures. For each template, we compare a JSON representation versus skipping the intermediate generation step (Rep), maintaining the same style parameters in both settings.
Experiments with additional models.
[…]
Figure 3: Reasons annotators preferred each document. While annotators largely preferred documents generated with an intermediate representation, the most common reasons for preference are better formatting and information content. We exclude the “Other” count as it was only selected once.
@@ 6.2 Human evaluation
After showing that LLMs benefit from intermediate structured representations in document transformations, we investigate whether our proposed evaluation framework aligns better with human judgment than previously proposed metrics. We sample 100 documents each from DOC2PPT and LongSumm, and use the entirety of the Paper-Poster dataset in this study. We present annotators with 2 versions of each document, one generated with the intermediate representation and one without. Both versions use gpt4-32k, as the best performing model.
The annotators are provided with the original paper and the intended document type (blog, slide deck, or poster), and are asked the following questions: (1) Which document do you prefer? (2) On a scale of 1-3, to what degree do you prefer your selection? (3) Why do you prefer your selection? For question 3, annotators are also provided with a multi-select checklist of reasons for their preference: (1) quality of the content, (2) formatting, (3) document style matching the intended document types, (4) information represented in the document, and (5) other (along with a free text box). The full instructions, including the reasons provided and examples, can be found in Appendix §B.
[…]
@@
The documents generated with an intermediate representation were preferred by 82%, based on majority vote (71% unanimously). The annotator agreement score was 0.51 with Krippendorff’s alpha, indicating that while this is a subjective and specialized task, even non-expert annotators agree to a moderate degree. A visualization of the reasons the annotators preferred their selection can be found in Figure 3. It can be seen that while annotators largely preferred the documents generated with an intermediate representation, the most common reasons for preference are better formatting and better information content. This indicates that the structure provided by the intermediate representation makes it easier for the model to format the final document well. Additionally, the intermediate representation only includes the most salient information from the original text, resulting in higher quality of information content. Finally, we see a fairly even distribution across different templatic views for the reasons of preference, indicating humans prefer the documents generated with the intermediate representation across different document types.
Which metric correlates better with humans?
We test whether our metric, as described in §5, correlates better with human preference compared to prior evaluation metrics in the literature. For each annotation, given the degree of the preference $d$ (Appendix §B Q. 2) we convert value to a score $P(d)\rightarrow[1,2,3]$ if $d$ is slight, moderate, or strong, respectively. If the annotator prefers the document generated without an intermediate representation, we take $-P(d)$ instead. This allows us to measure if the metric captures directionality of preference along with degree. in parallel, we compute the automatic score $m$ for each metric, then calculate $S=m(\textrm{with rep})-m(\textrm{skip rep})$ where $m$ is the metric we are evaluating (e.g ROUGE). If a human annotator prefers a document generated without the intermediate step, we’d expect a good metric to assign a higher score to that document as well, resulting in both $S$ and $P(d)$ being negative (and positive in the opposite case). Using this intuition we assign an affinity score of a metric with respect to human evaluation as the Pearson correlation Freedman et al. (2007) of $S$ and $P(d)$.
Since prior metrics are not designed to account for the structure of documents, we compute them by extracting only the text of both the generated and reference documents. The correlations with human judgement for each metric to its respective TAE variants can be found in Table 3. As we can see from the results, evaluations using our template adaptable framework correlate more highly with human judgement, except in the case of BERTScore. In the latter case the results are not statistically significant, and we hypothesize that the open-domain nature of BERT embeddings are poorly suited to represent the semantic similarity of scientific text.
@@
TAE BLEU 13.8
[…]
TAE BERTScore 5.4*
Table 3: Correlation of evaluation metrics with human judgement. We compare each metric computed using the TAE framework versus the standard computation. *Indicates the correlation is not statistically significant ($p>0.01$).
@@ 7 Conclusion
In many domains, people choose to disseminate information across different modalities and formats for better communication to broader audiences. We proposed a unified view of document transformation and evaluation. We showed that LLMs are capable of templatic document generation with minimal supervision, and that a structured, intermediate representation can improve performance, particularly for smaller models. We also introduced a flexible precision-recall framework for automatic evaluation that easily integrates existing evaluation metrics into a unified system and allows for comparison across diverse datasets without additional task specific metric design. Finally, we conducted a human evaluation and showed that annotators prefer the documents generated using the intermediate representation 82% of the time and that our evaluation framework correlates better with human preference than standard evaluation metrics.
8 Limitations
Although our methods are not domain specific, we only evaluated them in the scientific domain, due to the availability of public data. Additionally, our framework is limited to textual content. In future work we plan to explore the application of our unified framework for generation and evaluation on document views in other domains, as well as incorporating multi-modal models and content generation. Finally, it is possible that some of our test data has leaked into the training data of the models with which we experimented. This limitation is not unique to our work and exists for our baselines in addition to our methods.
9 Ethics
The potential risks of our work are similar to those of other work in downstream applications of LLMs. LLM generated documents can potential generate copy-righted material Carlini et al. (2020), personally-identifiable information Lukas et al. (2023), or factually incorrect text Manakul et al. (2023). The use of LLMs to generate documents may violate some academic dishonesty policies Zdravkova et al. (2023). Our system is intended to be used in collaboration with human writers. Users should edit the generations, checking for factual inconsistencies and other potential errors. Our work is intended to save users time that might be spent repeating information across multiple documents, so they can focus on content creation. Therefore, we believe the benefits of our work outweigh the potential risks.
Appendix A Prompt details
The prompts and intermediate representation template used can be found in Table 4 and Figure 4, respectively. We note that the specific structure provided to the prompt is not inherent to our method, and a different structure could be provided depending on the input document and domain. For the tasks we evaluate in this paper, we use the following style parameters:
  • Blogs: “Blogs should include paragraphs introducing the topic, a summary of the input document, and important takeaways. The blog should be more readable to a general audience than the input document.”
Figure 4: Template of the intermediate representation provided to the prompts in Table 4.
Appendix B Annotation Instructions
B.1 Questions